Can you please help me with this proportion? The ratio of the measures of three angles of a triangle is 5:7:8. Find the measure of each angle of the triangle.
Answer by checkley71(8403)

5X+7X+8X=180
20X=180
X=180/20
X=9 THUS THE ANGLES ARE
5*9=45
7*9=63
8*9=72
PROOF
45+63+72=180
180=180